One canned juice drink is 25 percent orange juice and the other is 10 percent juice How many Liters of each should be mixed together in order to get 15Liters that is 17 percent orange juice?
Answer by checkley77(12844)

.25x+.10(15-x)=.17*15
.25x+1.5-.10x=2.55
.15x=2.55-1.5
.15x=1.05
x=1.05/.15
x=7 L of 25% oj is used.
15-7=8 L of .10% oj is used.
Proof:
.25*7+.10*8=2.55
1.75+.8=2.55
2.55=2.55